Jamie bought $15 worth of drinks from Starbucks. Lattes cost $3 and coffees cost $1.50. Jamie purchased 6 drinks. How many of each type of drink did she buy?
Answer by macston(5194)

L=lattes; C=coffees
.
L+C=6
L=6-C
.
$3L+$1.50C=$15
$3(6-C)+$150C=$15
$18-$3C+$1.50C=$15
-$1.50C=-$3
C=2
ANSWER 1: She bought 2 coffees.
.
L=6-C=6-2=4
ANSWER 2: She bought 4 lattes.
.
CHECK:
$3.00L+$1.50C=$15.00
$3.00(4)+$1.50(2)=$15.00
$12.00+$3.00=$15.00
$15.00=$15.00
.